# Leray-Schauder Mappings for Operator Learning

### v2

@@ 1. Introduction
Operator learning is a branch of deep learning involved with approximating (potentially highly nonlinear) continuous operators between Banach spaces. The interest of operator learning lies in the fact that it allows to model complex phenomena, e.g. dynamical systems, whose underlying governing equations are not known [DeepOnet, ANIE, NIDE]. The study of operator learning was initiated by the theoretical work [chen], whose implementation was given in [DeepOnet].
@@
The problem of operator learning is therefore to model mappings between infinite dimensional spaces. Since most of the algorithms in practice discretize the domains of the functions, and therefore create a finite space upon which the neural networks are used, one might ask whether the mapping is really between function spaces, or it can just be reduced to some map between high dimensional spaces. For instance, one might ask whether it is possible to upsample the discretized domains after training, e.g. to interpolate and output arbitrarily dense predictions, which would show that the model does not depend on the spacetime stamps used during training. In other words, to what extent the model operates in the infinite dimensional spaces, rather than on some high dimensional Euclidean spaces obtained via discretization.
@@
Various models in operator learning deal with such a question in different ways. For intance, [DeepOnet] fixes points in the domain, but allows continuous inputs. The articles [NODE, ANIE, NIDE] use solvers (for ODEs, IEs or IDEs) to allow continuous sampling/output of points. In [FNO], the use of Fourier transforms allows to upsample the domain of space of functions.
@@
Our perspective in the present work stems from the idea of learning an operator between bases of functions as in [Spectral], where the Chebyshev polynomials have been used for projections as in the Galerkin methods [Fle]. In fact, it was proved in [EZ_proj] that using Leray-Schauder mappings to “nonlinearly project” over a basis of chosen elements approximating a compact set of a Banach space, one can approximate any (possibly nonlinear) operator. This addresses the aforementioned question of learning operators between infinite dimensional spaces rather than just discretization of them, because the approximation happens at the functional level.
@@
In fact, in the present article we introduce a family of neural networks that approximate a compact subseet of a Banach space obtained via Leray-Schauder mappings. Then, the operator learning problem is formulated between the coordinate spaces of the neural networks. This allows us to treat the problem without any discretization, but by using the finite-dimensional reduction provided by the Leray-Schauder mappings to approximate the original learning problem. We show that there exist theoretical guarantees for the approximation to be arbitrarily accurate. In practice, therefore, in this article we also learn the elements upon which the Leray-Schauder mappings project the compact subspace. This represents a paradigmatic shift with respect to the setting of [Spectral], where the projection used is fixed.
@@
We numerically demonstrate the efficacy of the model, and show that it obtains results comparable to state of the art models. The tasks used for comparison are a dataset of integral equations, which shows the efficacy of the model on a task of learning nonlocal operators, and a dataset of Burger’s equation, which explores a PDE problem. The codes for the implementation of this work are found at https://github.com/emazap7/Leray_Schauder_neural_net.
@@ 2. Theoretical Preliminaries
Let $X$ and $Y$ be Banach spaces, let $T:X\longrightarrow Y$ be a continuous (possibly nonlinear) map, and let $K\subset X$ be a compact subset. Then, for any choice of $\epsilon>0$ there exist natural numbers $n,m\in\mathbb{N}$, finite dimensional subspaces $E_{n}\subset X$ and $E_{m}\subset Y$, continuous maps $P_{n}:X\longrightarrow E_{n}$ and $P_{m}:Y\longrightarrow E_{m}$, and a neural network $f_{n,m}:\mathbb{R}^{n}\longrightarrow\mathbb{R}^{m}$ such that for every $x\in K$
(1) $||T(x)-\phi_{m}^{-1}f_{n,m}\phi_{n}P_{n}(x)||<\epsilon,$
@@
for all $i=1,\ldots,n$.
We will refer to these operators as Leray-Schauder projections, although they are not linear, following the same convention as [Topological], which was already adopted in [EZ_proj].
@@
Let $X$ be the Banach space of continuous functions over a compact $K\subset\mathbb{R}^{N}$ with the uniform norm, $C(K,\mathbb{R}^{M})$. Let $T:X\longrightarrow X$ be a continuous operator, and let $\mathbb{K}\subset X$ be a compact subset. Then, it is possible to find neural networks $\{g_{i}:\mathbb{R}^{N}\longrightarrow\mathbb{R}^{M}\}_{i=1}^{n}$ and $f_{n}:\mathbb{R}^{n}\longrightarrow\mathbb{R}^{n}$ such that for every $x\in\mathbb{K}$
(4) $||T(x)-\phi_{n}^{-1}f_{n}\phi_{n}\tilde{P}_{n}(x)||_{\infty}<\epsilon,$
where $\phi_{k}:F_{k}\longrightarrow\mathbb{R}^{k}$ indicates an isomorphism between the finite dimensional space $F_{k}$ and $\mathbb{R}^{k}$, and $\tilde{P}_{n}$ indicates the Leray-Schauder mapping on the space spanned by the neural networks $g_{i}$.
@@
We fix $\epsilon>0$. Using the Heine-Cantor theorem and the construction of Leray-Schauder mappings, it is possible to find functions $x_{i}$, with $i=1,\ldots,n$, such that
(5) $\|x-P_{n}(x)\|_{\infty}$ $<$ $\frac{\epsilon}{6},$
@@
for all $x,z_{1},z_{2}\in\mathbb{K}$. Here the continuous maps $P_{n}$ map $\mathbb{K}$ in the space $E_{n}$ spanned by $x_{i}$, $i=1,\ldots,n$.
The fundamental step of the present proof is to show that we can find neural networks $g_{i}$ spanning a space $F_{n}$ which has the property that $\|x-\tilde{P}_{n}x\|_{\infty}<\frac{\epsilon}{3}$ for all $x\in\mathbb{K}$, where $\tilde{P}_{n}$ is a mapping on $F_{n}$ which we define below. This would allow us to replace the approximation in (5), by an analogous approximation using mappings $\tilde{P}$ and a neural network basis instead of the functions $x_{i}$.
@@
$\tilde{\mu}_{i}(x)$ $=$ $\begin{cases}\epsilon-\|x-g_{i}\|_{\infty},\hskip 28.45274pt\|x-g_{i}\|_{\infty}\leq\epsilon\\ 0,\hskip 89.626pt\|x-g_{i}\|_{\infty}>\epsilon.\end{cases}$
In addition, upon taking $g_{i}$ such that $\|x_{i}-g_{i}\|_{\infty}$ is sufficiently small, we can make $|\mu_{i}(x)-\tilde{\mu}_{i}(x)|$ arbitrarily small as well. We prove that
(8) $\inf_{x\in\mathbb{K}}\sum_{j=1}^{n}\mu_{j}(x)>0.$
By way of contradiction, assume otherwise. Then, let $\{y_{k}\}$ be a sequence in $\mathbb{K}$ such that as $k\longrightarrow+\infty$ we have $\sum_{j=1}^{n}\mu_{j}(y_{k})\longrightarrow 0$. Since $\mathbb{K}$ is a compact in a metric space, it is sequentially compact. Let $y_{k_{r}}$ be a convergent subsequence, and let $y_{k_{r}}\longrightarrow y^{*}$. The set $\{x_{1},\ldots,x_{n}\}$ is an $\epsilon$-net by construction, therefore we can find $x_{i}$ such that $q:=\|y^{*}-x_{i}\|_{\infty}<\epsilon$, hence $\epsilon-q>0$. Let $0<\delta<\epsilon-q$. By continuity of the map $\sum_{j=1}^{n}\mu_{i}:\mathbb{K}\longrightarrow\mathbb{R}^{+}_{0}$ we can find a $\delta>0$ such that for all $x\in B(y^{*},\delta)\bigcap\mathbb{K}$ we have $\sum_{j=1}^{n}\mu_{j}(x)\geq\epsilon-q-\delta>0$. Therefore, for $r$ large enough, we have $\sum_{j=1}^{n}\mu_{j}(y_{k_{r}})\geq\epsilon-q-\delta>0$ against the fact that $\sum_{j=1}^{n}\mu_{j}(y_{k})\longrightarrow 0$. This shows that (8) holds true. We put $m:=\inf_{x\in\mathbb{K}}\sum_{j=1}^{n}\mu_{j}(x)$.
Suppose $|\mu_{i}(x)-\tilde{\mu}_{i}(x)|<q$, where $q$ is small enough such that $\sum_{i=1}^{n}\mu_{i}(x)-q>0$. This can be done since $m>0$ as shown above, and it guarantees that $\sum_{j=1}^{n}\tilde{\mu}_{j}(x))$ does not vanish. Then, for all $i=1,\ldots,n$, we have that
@@
$\leq$ $\sum_{j=1}^{n}\frac{1}{m(m-q)}[|\mu_{i}(x)\tilde{\mu}_{j}(x)-\mu_{i}(x)\mu_{j}(x)|+|\mu_{i}(x)\mu_{j}(x)-\tilde{\mu}_{i}(x)\mu_{j}(x)|]$
$=$ $\sum_{j=1}^{n}\frac{1}{m(m-q)}[|\mu_{j}(x)|\cdot|\tilde{\mu}_{j}(x)-\mu_{j}(x)|+|\mu_{j}(x)|\cdot|\mu_{i}(x)-\tilde{\mu}_{i}(x)|]$
$<$ $\frac{2n\epsilon q}{m(m-q)},$
which implies that $|\frac{\mu_{i}(x)}{\sum_{j=1}^{n}\mu_{j}(x)}-\frac{\tilde{\mu}_{i}(x)}{\sum_{j=1}^{n}\tilde{\mu}_{j}(x)}|\longrightarrow 0$ as $q\longrightarrow 0$. Since
$|\sum_{i=1}^{n}\frac{\mu_{i}(x)}{\sum_{j=1}^{n}\mu_{j}(x)}-\sum_{i=1}^{n}\frac{\tilde{\mu}_{i}(x)}{\sum_{j=1}^{n}\tilde{\mu}_{j}(x)}|$$\leq$ $\sum_{i=1}^{n}|\frac{\mu_{i}(x)}{\sum_{j=1}^{n}\mu_{j}(x)}-\frac{\tilde{\mu}_{i}(x)}{\sum_{j=1}^{n}\tilde{\mu}_{j}(x)}|$
$<$ $\frac{2n^{2}\epsilon q}{m(m-q)},$
we find that by taking $g_{i}$ such that $\|x_{i}-g_{i}\|_{\infty}$ is small enough, we can obtain $|\sum_{i=1}^{n}\frac{\mu_{i}(x)}{\sum_{j=1}^{n}\mu_{j}(x)}-\sum_{i=1}^{n}\frac{\tilde{\mu}_{i}(x)}{\sum_{j=1}^{n}\tilde{\mu}_{j}(x)}|<\frac{\epsilon}{12\max_{i}\|g_{i}\|_{\infty}}$ and at the same time have $\|x_{i}-g_{i}\|_{\infty}<\frac{\epsilon}{12}$.
@@
$\|x-\tilde{P}_{n}x\|_{\infty}$
$<$ $\|x-P_{n}x\|_{\infty}+\|P_{n}x-\tilde{P}_{n}x\|_{\infty}$
$=$ $\frac{\epsilon}{6}+\|\frac{\sum_{i}\mu_{i}(x)x_{i}}{\sum_{j}\mu_{j}(x)}-\frac{\sum_{i}\tilde{\mu}_{i}(x)g_{i}}{\sum_{j}\tilde{\mu}_{j}(x)}\|_{\infty}$
$\leq$ $\frac{\epsilon}{6}+\sum_{i=1}^{n}\|\frac{\mu_{i}(x)x_{i}}{\sum_{j=1}^{n}\mu_{j}(x)}-\frac{\tilde{\mu}_{i}(x)g_{i}}{\sum_{j=1}^{n}\tilde{\mu}_{j}(x)}\|_{\infty}$
$\leq$ $\frac{\epsilon}{6}+\sum_{i=1}^{n}[\|\frac{\mu_{i}(x)x_{i}}{\sum_{j=1}^{n}\mu_{j}(x)}-\frac{\mu_{i}(x)g_{i}}{\sum_{j=1}^{n}\mu_{j}(x)}\|_{\infty}+\|\frac{\mu_{i}(x)g_{i}}{\sum_{j=1}^{n}\mu_{j}(x)}-\frac{\tilde{\mu}_{i}(x)g_{i}}{\sum_{j=1}^{n}\tilde{\mu}_{j}(x)}\|_{\infty}]$
$\leq$ $\frac{\epsilon}{6}+\sum_{i=1}^{n}[\frac{\mu_{i}(x)}{\sum_{j=1}^{n}\mu_{j}(x)}\|x_{i}-g_{i}\|_{\infty}+\|g_{i}\|_{\infty}|\frac{\mu_{i}(x)}{\sum_{j=1}^{n}\mu_{j}(x)}-\frac{\tilde{\mu}_{i}(x)}{\sum_{j=1}^{n}\tilde{\mu}_{j}(x)}|]$
$<$ $\frac{\epsilon}{6}+\frac{\epsilon}{12}+\frac{\epsilon}{12}$
$=$ $\frac{\epsilon}{3},$
@@
where we have used the fact that $\mu_{j}(x)\geq 0$ by definition of Leray-Schauder mapping. A similar procedure can be used to find mappings $\tilde{P}_{n^{\prime}}$ for the compact $T(\mathbb{K})$. However, upon possibly passing to subrefinements, we can assume that $n=n^{\prime}$ and $\tilde{P}_{n^{\prime}}=\tilde{P}_{n}$. This applies in the particularly important case when $T(\mathbb{K})\subset\mathbb{K}$.
@@
We let $\phi_{n}$ be an isomorphism which identifies $F_{n}$ with $\mathbb{R}^{n}$, and construct the mapping $T_{n}$ induced by $T$ on $F_{n}$. We find a neural network architecture which approximates $T_{n}$ on the compact $\phi_{n}P_{n}(\mathbb{K})$. With these definitions one can verify that the approximation (4) holds for all $x\in\mathbb{K}$, by adding a bound of type $\|(\phi_{n}^{-1}f_{n}\phi_{n})P_{n}(x)-(\phi_{n}^{-1}f_{n}\phi_{n})\tilde{P}_{n}(x)\|_{\infty}<\epsilon$, using the uniform continuity of $f_{n}$ on the compact $\phi_{n}P_{n}(\mathbb{K})\bigcup\phi_{n}\tilde{P}_{n}(\mathbb{K})$. ∎
@@
The construction of $\tilde{P}_{n}$ is complicated by the fact that we need to map on a space spanned by neural networks through the use of the functions $\tilde{\mu}_{i}$, where we need to find $g_{i}$ such that the maps $\tilde{\mu}_{i}$ have the required properties. This presents a significant implementational, and practical, problem in terms of constructing the mappings $\tilde{P}_{n}$. In practice, we would like to implement $\tilde{P}_{n}$ as neural networks as well, so that the whole architecture consists of neural network components. In the rest of this section we will work toward this direction, showing that it is indeed possible to replace $\tilde{P}_{n}$ by neural networks. This construction will be seen to be a universal approximation architecture on functions that are regular enough in the sense of Hölder. The result is therefore somewhat weaker than the case when one uses $\tilde{P}_{n}$ maps as above. We will denote by $L^{P}_{\mu}([a,b])$ the $L^{p}$-space on $[a,b]$ for some measure $\mu$ (which is assumed to be sufficiently regular). In what follows, we denote the Hölder space by $C^{\kappa,\alpha}([a,b])$, where $\kappa\geq 1$, and $\alpha\in(0,1]$. We assume that $p>2\kappa-1$. We will later mention how to generalize the result to the multivariate case. This relies on cubature formulas, and our choice of proceeding with the single varaible case is motivated by the wish of keeping the notation as simple as possible.
@@
For an interval $[a,b]$ of $\mathbb{R}$, for a discretization (e.g. a partition) $\{t_{1},\ldots,t_{k}\}\subset[a,b]$ and a neural network $F:\mathbb{R}^{k}\longrightarrow\mathbb{R}^{s}$, we define $F(f):=F(f(t_{1}),\ldots,f(t_{k}))$ for any function $f:[a,b]\longrightarrow\mathbb{R}$, and the components of $F$ in $\mathbb{R}^{s}$ are indicated by $F^{i}(f)$. In other words, $F$ is applied to the evaluation of $f$ on the nodes of discretization. In machine learning, one usually has access to data sampled on discretizations of the domains of definition. Therefore, the neural network can work as a proxy of an operator. We leverage this fact to approximate the Leray-Schauder mappings $P_{n}$ via neural networks.
@@
Let $\mathbb{K}\subset L^{p}_{\mu}([a,b])$ be a compact subspace, and let $P_{n}$ be the Leray-Schauder map $P_{n}:\mathbb{K}\longrightarrow E_{n}$, where $E_{n}$ is spanned by an $\epsilon$-net of smooth functions $\{x_{1},\ldots,x_{n}\}$. Then, for any $\eta>0$, there exists a discretization $\{t_{1},\ldots,t_{k}\}$ of $[a,b]$, and a neural network $F:\mathbb{R}^{\times k}\longrightarrow\mathbb{[}0,+\infty)$ such that
@@
for all $x\in\mathbb{K}\bigcap\bar{B}_{\rho}$, where $B_{\rho}$ is an open ball in the Hölder sapce $C^{\kappa,\alpha}([a,b])$. for any $\rho>0$ such that $\max_{i}\|x_{i}\|_{\infty}\leq\rho$.

### v3

@@ 1. Introduction
Operator learning is a branch of deep learning involved with approximating (potentially highly nonlinear) continuous operators between Banach spaces. The interest of operator learning lies in the fact that it allows to model complex phenomena, e.g. dynamical systems, whose underlying governing equations are not known [DeepOnet, ANIE, NIDE]. The study of operator learning was initiated by the theoretical work [chen], whose implementation was given in [DeepOnet], and whose error bounds were further developed in [LMK].
@@
The problem of operator learning is therefore to model mappings between infinite dimensional spaces. Since most of the algorithms in practice discretize the domains of the functions, and therefore create a finite space upon which the neural networks are used, one might ask whether the mapping is really between function spaces, or it can just be reduced to some map between high dimensional spaces. For instance, one might ask whether it is possible to upsample the discretized domains after training, e.g. to interpolate and output arbitrarily dense predictions, which would show that the model does not depend on the spacetime stamps used during training, performing an interpolation task. In other words, to what extent the model operates in the infinite dimensional spaces, rather than on some high dimensional Euclidean space obtained via discretization. An example where learning on discretized spaces shows its limitations was pointed out in [cont_transformer], where the tokenization process in the transformer architecture resulted in the model learning step functions. In other words, the discretization of the input resulted in the discretization of the output, and poor interpolation results during test.
@@
Various models in operator learning deal with such a question in different ways. For intance, [DeepOnet] fixes points in the domain, but allows continuous inputs. The articles [NODE, ANIE, NIDE] use solvers (for ODEs, IEs or IDEs) to allow continuous sampling/output of points. In [FNO], the use of Fourier transforms allows to upsample the domain of space of functions. In [cont_transformer], the use of Sobolev-type norms to limit sharp edges due to steep derivatives is employed to regularize the step-like behavior induced by tokenization.
@@
Our perspective in the present work stems from the idea of learning an operator between bases of functions as in [Spectral], where the Chebyshev polynomials have been used for projections as in the Galerkin methods [Fle]. In fact, it was proved in [EZ_proj] that using Leray-Schauder mappings to “nonlinearly project” over a basis of chosen elements approximating a compact set of a Banach space, one can approximate any (possibly nonlinear) continuous operator. This addresses the aforementioned question of learning operators between infinite dimensional spaces rather than just discretization of them, because the approximation happens at the functional level.
@@
In fact, in the present article we introduce a family of neural networks that approximate a compact subset of a Banach space obtained via Leray-Schauder mappings. Then, the operator learning problem is formulated between the coordinate spaces of the neural networks. This allows us to treat the problem without any discretization, but by using the finite-dimensional reduction provided by the Leray-Schauder mappings to approximate the original learning problem. We show that there exist theoretical guarantees for the approximation to be arbitrarily accurate. In practice, therefore, in this article we also learn the elements upon which the Leray-Schauder mappings project the compact subspace. This represents a paradigmatic shift with respect to the setting of [Spectral], where the projection used is fixed.
@@
We numerically demonstrate the efficacy of the model, and show that it obtains results comparable to state of the art models. We test the model on interpolation tasks, to demonstrate the independence of the model on the grid size used during training. The systems used for comparison consist of a dataset of integral equations, which shows the efficacy of the model on a task of learning nonlocal operators, and a dataset of Burger’s equation, which explores a PDE problem. The codes for the implementation of this work are found at https://github.com/emazap7/Leray_Schauder_neural_net.
@@ 2. Theoretical Construction
Let $X$ and $Y$ be Banach spaces, let $T:X\longrightarrow Y$ be a continuous (possibly nonlinear) map, and let $K\subset X$ be a compact subset. Then, for any choice of $\epsilon>0$ there exist natural numbers $n,m\in\mathbb{N}$, finite dimensional subspaces $E_{n}\subset X$ and $E_{m}\subset Y$, a continuous map $P_{n}:X\longrightarrow E_{n}$, and a neural network $f_{n,m}:\mathbb{R}^{n}\longrightarrow\mathbb{R}^{m}$ such that for every $x\in K$
[…]
@@
for all $i=1,\ldots,n$. We observe that we allow the case where $x_{i}$ might not be in $K$.
We will refer to these operators as Leray-Schauder maps, or Leray-Schauder projections, although they are not linear. As shown in [EZ_proj], one can choose the elements $x_{i}$ in such a way that $P_{n}(x_{i})=x_{i}$, which would resemble a projection.
@@
In this article we will consider neural networks with continuous nonlinearities, so that the resulting neural networks are automatically continuous.
@@
Let $X$ and $Y$ be the Banach spaces of continuous functions over compacts $K_{1}\subset\mathbb{R}^{N_{1}}$ and $K_{2}\subset\mathbb{R}^{N_{2}}$ with the uniform norm, $X=C(K_{1},\mathbb{R}^{M_{1}})$ and $Y=C(K_{2},\mathbb{R}^{M_{2}})$. Let $T:X\longrightarrow Y$ be a continuous operator, and let $\mathbb{K}\subset X$ be a compact subset. Then, for a given arbitrary $\epsilon>0$, it is possible to find neural networks $\{g_{i}:\mathbb{R}^{N_{1}}\longrightarrow\mathbb{R}^{M_{1}}\}_{i=1}^{n}$, $\{h_{i}:\mathbb{R}^{N_{2}}\longrightarrow\mathbb{R}^{M_{2}}\}_{i=1}^{m}$, and $f_{n,m}$ such that for every $x\in\mathbb{K}$
[…]
where $\phi_{k}$ indicates an isomorphism between the finite dimensional space $F_{k}$ and the Euclidean space, and $\tilde{P}_{n}$ indicates the Leray-Schauder mapping on the space spanned by the neural networks $g_{i}$, and the $h_{i}$ neural networks span the space $E_{m}$.
@@
We fix $\epsilon>0$. As in [EZ_proj], we can split the proof in two steps, one where $T$ is uniformly continuous on a neighborhood $U$ of $\mathbb{K}$, and a general step where the initial map is approximated by a map uniformly continuous on a neighborhood $U$ of $\mathbb{K}$. Since the latter step is independent of the current setup, we can assume that $T$ is uniformly continuous on $U\supset\mathbb{K}$. It is possible to find functions $x_{i}$, with $i=1,\ldots,n^{\prime}$, such that
(5) $\|x-P_{n}(x)\|_{\infty}$ $<$ $\frac{\delta}{3},$
@@
for all $x\in\mathbb{K}$, and for all $z_{1},z_{2}\in U$ with $\|z_{1}-z_{2}\|_{\infty}<\delta$. Moreover, $\delta$ is chosen such that $B(x,\delta)\subset U$ for each $x\in\mathbb{K}$. Here the continuous maps $P_{n}$ map $\mathbb{K}$ in the space $E_{n}$ spanned by $x_{i}$, $i=1,\ldots,n^{\prime}$, where $\dim E_{n}=n$. In what follows we assume that $n^{\prime}=n$ for simplicity of notation, as the proof would not change if otherwise.
The fundamental step of the present proof is to show that we can find neural networks $g_{i}$ spanning a space $F_{n}$ which has the property that $\|x-\tilde{P}_{n}x\|_{\infty}<\delta$ for all $x\in\mathbb{K}$, where $\tilde{P}_{n}$ is a mapping on $F_{n}$ which we define below. This would allow us to replace the approximation in (5), by an analogous approximation using mappings $\tilde{P}$ and a neural network basis instead of the functions $x_{i}$.
@@
$\tilde{\mu}_{i}(x)$ $=$ $\begin{cases}\frac{\delta}{3}-\|x-g_{i}\|_{\infty},\hskip 28.45274pt\|x-g_{i}\|_{\infty}\leq\frac{\delta}{3}\\ 0,\hskip 89.626pt\|x-g_{i}\|_{\infty}>\frac{\delta}{3}.\end{cases}$
In addition, upon taking $g_{i}$ such that $\|x_{i}-g_{i}\|_{\infty}$ is sufficiently small, we can make $|\mu_{i}(x)-\tilde{\mu}_{i}(x)|$ arbitrarily small as well. We have that
(8) $\min_{x\in\mathbb{K}}\sum_{j=1}^{n}\mu_{j}(x)>0,$
as a consequence of the continuity of the map $\sum_{j=1}^{n}\mu_{i}:\mathbb{K}\longrightarrow\mathbb{R}^{+}_{0}$, the compactness of $\mathbb{K}$, and the fact that for each $x\in\mathbb{K}$ there exists $i$ such that $\|x-x_{i}\|_{\infty}<\frac{\delta}{3}$, and therefore for which $\mu_{i}(x)>0$. We put $m_{0}:=\min_{x\in\mathbb{K}}\sum_{j=1}^{n}\mu_{j}(x)$.
Suppose $|\mu_{i}(x)-\tilde{\mu}_{i}(x)|<q$, where $q$ is small enough such that $\sum_{i=1}^{n}\mu_{i}(x)-q>0$. This can be done since $m>0$, and it guarantees that $\sum_{j=1}^{n}\tilde{\mu}_{j}(x)$ does not vanish. Then, for all $i=1,\ldots,n$, we have that
@@
$\leq$ $\sum_{j=1}^{n}\frac{1}{m_{0}(m_{0}-q)}[|\mu_{i}(x)\tilde{\mu}_{j}(x)-\mu_{i}(x)\mu_{j}(x)|+|\mu_{i}(x)\mu_{j}(x)-\tilde{\mu}_{i}(x)\mu_{j}(x)|]$
$=$ $\sum_{j=1}^{n}\frac{1}{m_{0}(m_{0}-q)}[|\mu_{j}(x)|\cdot|\tilde{\mu}_{j}(x)-\mu_{j}(x)|+|\mu_{j}(x)|\cdot|\mu_{i}(x)-\tilde{\mu}_{i}(x)|]$
$<$ $\frac{2n\delta q}{3m_{0}(m_{0}-q)},$
which implies that $|\frac{\mu_{i}(x)}{\sum_{j=1}^{n}\mu_{j}(x)}-\frac{\tilde{\mu}_{i}(x)}{\sum_{j=1}^{n}\tilde{\mu}_{j}(x)}|\longrightarrow 0$ as $q\longrightarrow 0$. We therefore find that by taking $g_{i}$ such that $\|x_{i}-g_{i}\|_{\infty}$ is small enough, we can obtain
$\sum_{i=1}^{n}|\frac{\mu_{i}(x)}{\sum_{j=1}^{n}\mu_{j}(x)}-\frac{\tilde{\mu}_{i}(x)}{\sum_{j=1}^{n}\tilde{\mu}_{j}(x)}|$ $<$ $\frac{\delta}{3\max_{i}\|g_{i}\|_{\infty}},$
and at the same time have $\|x_{i}-g_{i}\|_{\infty}<\frac{\delta}{3}$.
@@
$\|x-\tilde{P}_{n}x\|_{\infty}$ $\leq$ $\|x-P_{n}x\|_{\infty}+\|P_{n}x-\tilde{P}_{n}x\|_{\infty}$
$=$ $\frac{\delta}{3}+\|\frac{\sum_{i}\mu_{i}(x)x_{i}}{\sum_{j}\mu_{j}(x)}-\frac{\sum_{i}\tilde{\mu}_{i}(x)g_{i}}{\sum_{j}\tilde{\mu}_{j}(x)}\|_{\infty}$
$\leq$ $\frac{\delta}{3}+\sum_{i=1}^{n}\|\frac{\mu_{i}(x)x_{i}}{\sum_{j=1}^{n}\mu_{j}(x)}-\frac{\tilde{\mu}_{i}(x)g_{i}}{\sum_{j=1}^{n}\tilde{\mu}_{j}(x)}\|_{\infty}$
$\leq$ $\frac{\delta}{3}+\sum_{i=1}^{n}[\|\frac{\mu_{i}(x)x_{i}}{\sum_{j=1}^{n}\mu_{j}(x)}-\frac{\mu_{i}(x)g_{i}}{\sum_{j=1}^{n}\mu_{j}(x)}\|_{\infty}+\|\frac{\mu_{i}(x)g_{i}}{\sum_{j=1}^{n}\mu_{j}(x)}-\frac{\tilde{\mu}_{i}(x)g_{i}}{\sum_{j=1}^{n}\tilde{\mu}_{j}(x)}\|_{\infty}]$
$\leq$ $\frac{\delta}{3}+\sum_{i=1}^{n}[\frac{\mu_{i}(x)}{\sum_{j=1}^{n}\mu_{j}(x)}\|x_{i}-g_{i}\|_{\infty}+\|g_{i}\|_{\infty}|\frac{\mu_{i}(x)}{\sum_{j=1}^{n}\mu_{j}(x)}-\frac{\tilde{\mu}_{i}(x)}{\sum_{j=1}^{n}\tilde{\mu}_{j}(x)}|]$
$<$ $\frac{\delta}{3}+\frac{\delta}{3}+\frac{\delta}{3}$
$=$ $\delta,$
@@
where we have used the fact that $\mu_{j}(x)\geq 0$ by definition of Leray-Schauder mapping. A similar procedure can be used to find mappings $\tilde{P}_{m}$ for the compact $T(\mathbb{K})$, where the centers of the $\frac{\epsilon}{3}$-balls in [EZ_proj] are taken to be neural networks.
@@
We let $\phi_{n}$ be an isomorphism which identifies $F_{n}$ with $\mathbb{R}^{\dim F_{n}}$, and construct the mapping $T_{n,m}$ induced by $T$ on $F_{n}$ as in Theorem 2.1 of [EZ_proj]. We find a neural network architecture which approximates $T_{n,m}$ on the compact $\phi_{n}\tilde{P}_{n}(\mathbb{K})$. With these definitions one can verify that the approximation (4) holds for all $x\in\mathbb{K}$, by considering the same chain of inequalities in Theorem 2.1 of [EZ_proj] with $\tilde{P}_{n}$ replacing $P_{n}$. ∎
Remark 2.3.
We note that the proof of Theorem 2.2 is showing an extra bit of information, other than the approximation of $T$. It is showing that this can be done by also approximating the analytically defined maps $P_{n}$ by neural network counterparts $\tilde{P}_{n}$.
Theorem 2.2 produces an approximation result similar to that of [chen], and consequently that of DeepONet [DeepOnet]. However, the present approach is based on significantly different methods. In particular, as a byproduct of this approach, the nonlinear maps used to “project” on a finite span of neural network basis functions are obtained by approximating Leray-Schauder maps whose analytical formulation is known a priori. These are the maps $\tilde{P}_{n}$ used in the proof above. The functionals used in DeepONet to nonlinearly “project” on the finite dimensional space of trunk networks, do not approximate analytically defined maps, and it is therefore not clear a priori what they need to approximate for the universal approximation results to hold. This fact has also implications on the algorithmic implementation given below, as the maps $\tilde{P}_{n}$ are implemented with neural networks explicitly.
@@
The construction of $\tilde{P}_{n}$, in practice, implies the fact that we need to map on a space spanned by neural networks through the use of the functions $\tilde{\mu}_{i}$, where we need to find $g_{i}$ such that the maps $\tilde{\mu}_{i}$ have the required properties. We would like to implement $\tilde{\mu}_{i}$ in $\tilde{P}_{n}$ as neural networks as well, so that the whole architecture consists of neural network components only. In the rest of this section we will work toward this direction, showing that it is indeed possible to replace $\tilde{P}_{n}$ by neural networks. This construction will be seen to be a universal approximation architecture on functions that are regular enough with respect to the norms of their derivatives. We will denote by $L^{p}_{\mu}(K)$ the $L^{p}$-space on a compact $K$ for the measure $\mu$ which is assumed to be Lebesgue and normalized on $K$ for simplicity, although the reasoning can be adapted to finite Borel measures. We consider $1\leq p<+\infty$, and we denote the Hölder space by $C^{\kappa,\alpha}([a,b])$, where $\kappa\geq 1$, and $\alpha\in(0,1]$. We assume that $p\geq\kappa+\alpha$. We will later mention how to generalize the result to the multivariate case. This relies on cubature formulas, and our choice of proceeding with the single varaible case first, is motivated by the wish of keeping the notation as simple as possible.
@@
For an interval $[a,b]$ of $\mathbb{R}$, for a discretization (e.g. a partition) $\{t_{1},\ldots,t_{k}\}\subset[a,b]$ and a neural network $F:\mathbb{R}^{k}\longrightarrow\mathbb{R}^{s}$, we define $F(f):=F(f(t_{1}),\ldots,f(t_{k}))$ for any function $f:[a,b]\longrightarrow\mathbb{R}$. In other words, $F$ is applied to the evaluation of $f$ on the nodes of discretization. In machine learning, one usually has access to data sampled on discretizations of the domains of definition. Therefore, the neural network can work as a proxy of an operator. We leverage this fact to approximate the Leray-Schauder mappings $P_{n}$ via neural networks. In what follows we restrict ourselves to quadrature rules that are continuous on their inputs. We will say that a function $u$ is sufficiently smooth if it is such that $|u|^{p}$ is in $C^{\kappa,\alpha}([a,b])$, e.g. it is $k+2$ times differentiable with all continuous derivatives in a neighborhood of $[a,b]$ since $p\geq\kappa+\alpha$.
@@
Let $\mathbb{K}\subset L^{p}_{\mu}([a,b])$ be a compact subspace, and let $P_{n}$ be the Leray-Schauder map $P_{n}:\mathbb{K}\longrightarrow E_{n}$, where $E_{n}$ is spanned by an $\epsilon$-net of sufficiently smooth functions $\{x_{1},\ldots,x_{n}\}$. Then, for any $\eta>0$, there exists a discretization $\{t_{1},\ldots,t_{k}\}$ of $[a,b]$, and neural networks $F^{i}:\mathbb{R}^{k}\longrightarrow\mathbb{R}$ such that
@@
for all $x\in\mathbb{K}\bigcap\bar{B}_{\rho}$, where $B_{\rho}$ is an open ball in the Hölder sapce $C^{\kappa+1,\alpha}([a,b])$, for any $\rho>0$ such that $\max_{i}\|x_{i}\|_{\infty}\leq\rho$.